OpenStudy (ingah):

Solve the following equation for x. logx+log(x-3)=log(3x) A: There are no solutions B: x= 0,6 C: x= 6 D: x= 0

rishavraj (rishavraj):

\[\log a + \log b = \log (ab)\]

OpenStudy (greatlife44):

\[\log_{10}x+\log_{10}(x-3) = \log_{10}3x \] we can apply the rules of logarithms. \[\log_{10}(x*(x-3)) = \log_{10}3x \] \[\log_{10}(x^{2}-3x) = \log_{10}(3x)\] \[\ln^{\log(x^{2}-3x)} = \ln^{\log{3x}}\] this allows us to simplify this expression into something simpler \[x^{2}-3x = 3x \]

OpenStudy (ingah):

Would it then be x=0? I see that 1 would work too, but that isn't an option

OpenStudy (greatlife44):

@ingah you can't take the log(0)

OpenStudy (greatlife44):

we were able to simplify this into an easier expression do you know how to solve this? \[x^{2}-3x = 3x \]

OpenStudy (ingah):

I tried plugging 6 in and it worked. (Earlier i mistakenly plugged in 36 instead of 6)

OpenStudy (greatlife44):

nice but let me show you how you can get to that

OpenStudy (greatlife44):

Please take a look at this @ingah \[x^{2}-3x-3x = 0\] \[x^{2}-6x = 0\] \[x(x-6) = 0 \] \[\frac{ x }{ x }(x-6) = \frac{ 0 }{ x }\] \[x-6 = 0 \] \[x =6 \]

OpenStudy (ingah):

Ohhhhhhh okay. Thank you!

OpenStudy (greatlife44):

yeah so used the properties of logarithms. in the previous example.